OpenStudy (mlgpro):

What is the value of the expression begin square root symbol a squared minus b plus nineteen end square root symbol minus b squared when a equals seven and b equals four? negative six negative seven negative eight three

OpenStudy (mlgpro):

@ImNotGoodAtMath

OpenStudy (tylerd):

\[\sqrt{7^2-4+10} - 4^2\]

OpenStudy (tylerd):

that look about right?

OpenStudy (tylerd):

instead of 10 19 typo

OpenStudy (mlgpro):

thanks

OpenStudy (tylerd):

its -8 but wanna know how to get there?

OpenStudy (mlgpro):

sure

OpenStudy (tylerd):

you cant move the 7^2 outside since theres numbers being added to it.

OpenStudy (tylerd):

7^2=49 49-4+19=64 sqrt of 64 = 8 4^2 = 16 8-16=-8

OpenStudy (anonymous):

tyler i recomend next time get hte person u are helping to do some of the work...

OpenStudy (anonymous):

and let him get the answer